OpenStudy (anonymous):

Please help Can the Values Be Added/Subtracted? 3 meters + 14 kilometers A. Yes, no converting needed B. Yes, but converting is required C. No, they cannot be added in any way

OpenStudy (gahm8684):

b

OpenStudy (ccswims):

^^^ what he said

OpenStudy (anonymous):

thanks

OpenStudy (gahm8684):

^^

OpenStudy (anonymous):

yeah, it's B <3

OpenStudy (anonymous):

Okay how about this one Solve for the Variable x + 5 = 16 15 - m = 4

OpenStudy (gahm8684):

x=11

OpenStudy (gahm8684):

and the other one is m=11

OpenStudy (anonymous):

thanks i think i'm good now :D

OpenStudy (gahm8684):

if you need any help tag me and close the question^-^

OpenStudy (gahm8684):

and ill help you

OpenStudy (anonymous):

x + 5 = 16 16 - 5 = 11 <3

OpenStudy (anonymous):

@gahm8684 lol im really needy today

OpenStudy (anonymous):

-13, -5, 14

OpenStudy (gahm8684):

yeah like that

OpenStudy (gahm8684):

my internet is very slow

OpenStudy (anonymous):

okee the last question is really easy so I'm just guna close the question thank you all for you help! :D <3

OpenStudy (anonymous):

no problem ! <3